OpenStudy (babyslapmafro):

The speed of light is 3.00x10^8m/s. How fast is this in mph? Use the conversion 1mile=1610m. Please show work.

OpenStudy (anonymous):

Are you familiar with doing a dimensional analysis?

OpenStudy (anonymous):

lets try an easier one first, lets convert 1 m/s into mph 1 m/s = 1*60 m/min = 60 m/min =60*60 m/hour = 3600 m/hour that's meters per hour, now we divide by that scale factor because it takes 1610m to make a whole mile 1 m/s = 3600/1610 mph = 2.2360248447... mph so now we can just multiply this number by the speed of light to get your answer. i make it about 670807453 mph

OpenStudy (babyslapmafro):

Thank you very much

OpenStudy (anonymous):

no problem :)